You stand atop a bridge 27.0 m above a river. You throw a rock horizontally off the bridge with a speed of 9.00 m/s. 

How much time elapses (in s) between the instant you release the rock and the instant it hits the water? 


1
Expert's answer
2021-02-08T18:39:59-0500
"h=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2h}{g}}=\sqrt{\dfrac{2\cdot27.0\ m}{9.8\ \dfrac{m}{s^2}}}=2.35\ s."